OpenStudy (jazzyfa30):

Solve for x if det = 24 | 2 x | |5 -3 |

OpenStudy (kirbykirby):

Do know this rule: \[\det\left[\begin{matrix}a & b \\c & d\end{matrix}\right]=ad-bc?\]

OpenStudy (jazzyfa30):

i've seen it before

OpenStudy (kirbykirby):

Ok so what we do is \[\det\left[\begin{matrix}2 & x \\ 5 & -3\end{matrix}\right]=2(-3)-(x)(5)=-6-5x\] But the problem tells us that \[\det \left[\begin{matrix}2 & x \\ 5 & -3\end{matrix}\right]=24\]

OpenStudy (kirbykirby):

Can you see what we can do next?

OpenStudy (jazzyfa30):

i think we multiply the det??? not sure

OpenStudy (kirbykirby):

No it's much easier than that. We have that the det is equal to -6-5x and equal to 24

OpenStudy (kirbykirby):

Thus you just need to solve -6-5x = 24, by simple algebra :)

OpenStudy (jazzyfa30):

oh so we add 6 to both sides then divide by 5

OpenStudy (kirbykirby):

yes, and then just multiply by (-1) since you'll have -x=___

OpenStudy (jazzyfa30):

-x=6 but dont we have to remove the -x which will turn it into x=-6

OpenStudy (kirbykirby):

yes, it's essentially the same operation :)

OpenStudy (jazzyfa30):

thanks do you mind helping me with another

OpenStudy (kirbykirby):

np. sure

OpenStudy (jazzyfa30):

Use the determinant to find the area of the triangle with the following vertices: (-3,1) (1,-3) (2,3) (remember area is positive, do not add units squared as the computer will count it incorrect.)

OpenStudy (kirbykirby):

I'm not sure why you need determinants. You can just use basic geometry to find the area of a triangle

OpenStudy (kirbykirby):

I've never used determinants to find the area of a triangle, but there is great info here: http://people.richland.edu/james/lecture/m116/matrices/applications.html

OpenStudy (jazzyfa30):

ive tried that site but i dont know where to gt my det from

OpenStudy (jazzyfa30):

@goformit100 help me please

OpenStudy (jazzyfa30):

hey @goformit100 do you know how to do the second question

OpenStudy (goformit100):

@ganeshie8

ganeshie8 (ganeshie8):

that site given by @kirbykirby explains the method. see if you can understand it... its bit tedious to explain :( you can come back and ask if u don't get something... give it a try :)

OpenStudy (jazzyfa30):

but where does the det come from

ganeshie8 (ganeshie8):

These are the vertices of triangle :- (-3,1) (1,-3) (2,3)

ganeshie8 (ganeshie8):

you need to fill them for 1st and 2nd columns

ganeshie8 (ganeshie8):

last column you fill wid 1's

ganeshie8 (ganeshie8):

-3 1 1 1 -3 1 2 3 1

ganeshie8 (ganeshie8):

find the determinant of this matrix

OpenStudy (jazzyfa30):

???

OpenStudy (kirbykirby):

how to find the determinant of a 3x3 matrix: http://www.youtube.com/watch?feature=player_embedded&v=21LWuY8i6Hw#at=271

OpenStudy (jazzyfa30):

det=-50

OpenStudy (jazzyfa30):

@ganeshie8 is that right

OpenStudy (jazzyfa30):

@GoldPhenoix i need some help |2 7 -3| |-1 5 -4| |6 9 0 |

OpenStudy (goldphenoix):

I don't know how to do matrices. Sorry.

OpenStudy (jazzyfa30):

kk

OpenStudy (jazzyfa30):

@Jhannybean please help

OpenStudy (jazzyfa30):

using the given. matrices find the value of its determinant: |2 7 -3| |-1 5 -4| |6 9 0 |

OpenStudy (jhannybean):

New problem?

OpenStudy (jazzyfa30):

no same one just put the hole problem up

OpenStudy (jhannybean):

Just follow the steps in the video stated by @kirbykirby . You'll be able to solve this :)

OpenStudy (jazzyfa30):

okay thanks